### v2

# Study of $1^{--}$ P wave charmoniumlike and bottomoniumlike tetraquark spectroscopy

Abstract
The masses of $1^{--}$ P-wave charmonium-like and bottomonium-like tetraquark states are calculated in a constituent quark model (CQM) where the Cornell-like potential and Breit-Fermi interaction are employed. All model parameters were imported from previous work, and predetermined by studying the low-lying conventional S- and P-wave light, charmed, bottom, charmonium, and bottomonium meson mass spectra. The decay widths of $1^{--}$ P-wave tetraquark states are calculated for possible two-body strong decay channels within the rearrangement mechanism, including $\omega\chi_{cJ}$, $\eta J/\psi$, and $\rho J/\psi$ for charmonium-like tetraquarks, and $\omega\chi_{bJ}$ for bottomonium-like tetraquarks. The tetraquark theoretical results are compared with the selected exotic states, also known as Y states, and tentative assignments are suggested. This study suggests that $\psi(4230)$, $\psi(4360)$, $\psi(4660)$, and $\Upsilon$(10753) may be P-wave tetraquark states and that multiple states might exist around 4.23 GeV and 4.36 GeV.
I Introduction
The study of exotic hadrons has gained significant attention in recent years, particularly in the context of the quark model. While the conventional meson and baryon spectra are well described by quantum chromodynamics (QCD) in terms of quark-antiquark and three-quark configurations, respectively, experimental discoveries of unconventional states challenge this simple classification. Among these exotic candidates, compact tetraquarks, bound states of two quarks and two antiquarks, offer a compelling framework to explain certain resonances that do not fit within the conventional meson spectrum Liu et al. (2019); Chen et al. (2023).
Charmonium-like tetraquarks, which contain a charm and anticharm quark pair along with an additional light quark pair, are particularly intriguing due to their potential connection to the so-called Y states observed in electron-positron annihilation experiments. In the latest naming scheme of the Particle Data Group (PDG) Navas et al. (2024), the Y states have been renamed as $\psi$ states because they share the same quantum number $1^{--}$, as the $J/\psi$ family.
[…]
The $\Upsilon(10753)$ bottomonium-like state was observed by the Belle Collaboration in the $e^{+}e^{-}\to\pi^{+}\pi^{-}\Upsilon(nS)$ process Mizuk et al. (2019). Recently, the Belle II Collaboration reported another observation that strongly supports the existence of the radiative transition process $\Upsilon(10753)\to\omega\chi_{bJ}(1P)$ Adachi et al. (2023). The energy dependence of the Born cross sections for $\Upsilon(10753)$ in the $\omega\chi_{bJ}(1P)$ channel Adachi et al. (2023) was found to be consistent with the shape of the $\Upsilon(10753)$ state in the $\pi^{+}\pi^{-}\Upsilon(nS)$ channel Mizuk et al. (2019). Belle-II concluded that the internal structure of $\Upsilon(10753)$ may differ from that of $\Upsilon(10860)$, where the latter is well understood as a predominantly conventional bottomonium state.
The exotic states exhibit properties that cannot be easily accommodated with conventional quarkonium expectations, suggesting the need for alternative interpretations. Plentiful theoretical pictures for charmonium-like exotic states, including charmonium core admixed with coupled channels Lu et al. (2017); Li and Chao (2009); Shah et al. (2012); Wang et al. (2019); Wang and Liu (2023), charmonium hybrids Liu et al. (2012); Luo and Liu (2006); Zhu (2005); Oncala and Soto (2017); Berwein et al. (2015), hidden-charm molecule statesChiu and Hsieh (2006); Qin et al. (2016); Ding (2009); Close and Downum (2009); Close et al. (2010), and hidden-charm tetraquark states Maiani et al. (2005); Ali et al. (2018); Drenska et al. (2009); Ebert et al. (2008); Wang (2024a, b, 2019a), have been employed to explore their internal structure and mass spectra, providing valuable insights into their nature.
Since the inception of the quark model Gell-Mann (1964), the tetraquark picture has been widely employed and further developed to understand the internal structure of exotic states Jaffe (1977a, b), not only in charmonium-like systems. For exotic states in the four-heavy-quark region, one may refer to the fully-heavy tetraquark picture Bai et al. (2025); Wu and Zhu (2025); Mistry and Majethiya (2024); Liu and Chen (2025); Bai et al. (2024); Wu et al. (2025); Chen et al. (2024); Wu et al. (2024); Zhang et al. (2024); Yu et al. (2023); Zhang et al. (2022); Mutuk (2022); Wang et al. (2022); Chen et al. (2022); Hu et al. (2023); Liu et al. (2021); Wang et al. (2021a, b). Similarly, for light exotic states, the fully-light tetraquark picture is considered Zhao et al. (2022); Agaev et al. (2019); Ebert et al. (2009); Wang et al. (2024); Xin and Wang (2024); Santopinto and Galata (2007); Lodha and Rai (2024)
In the previous work, the Y(4230) and the $\Upsilon(10753)$ cannot be accommodated in a conventional heavy quarkonium picture including S-D mixing Zhao et al. (2024). In this work, we focus on studying the mass spectrum and decay properties of the $1^{--}$ P-wave charmonium-like and bottomonium-like compact tetraquark states, referred to $Y_{c}$ and $Y_{b}$ states for convenience, within a constituent quark model framework.
The paper is organized as follows. In Sec. II, all the possible configurations of color, spin, and spatial degrees of freedom of tetraquark states are introduced, and a constituent quark model developed from our previous work Zhao et al. (2021, 2022) is briefly reviewed. In Sec. III, $1^{--}$ P-wave tetraquark mass spectra and decay branching ratios are evaluated in the constituent quark model. The theoretical results are compared with experimental data of Y states and tentative assignments for $1^{--}$ tetraquarks are suggested. A summary is given in Sec. IV. The details of the tetraquark spatial wave function, and cross term between different configuration from one-gluon exchange interaction are shown in the Appendix.
II THEORETICAL MODEL
II.1 Quark configurations of tetraquark
The construction of tetraquark states follows to the principle that a tetraquark state must be a color singlet, meaning the color wave function of the tetraquark must be a $[222]_{1}$ singlet within the $SU_{c}(3)$ group. In this work we consider only charmonium-like and bottomonium-like ($q_{1}Q_{2}\bar{q}_{3}\bar{Q}_{4}$) states, where $q_{1}$ and $Q_{2}$ are light quark and heavy quark, and $\bar{q}_{3}$ and $\bar{Q}_{4}$ are light antiquark and heavy antiquark respectively.
The permutation symmetry of the two-quark cluster ($qQ$) in tetraquark states is described by the Young tabloids $[2]_{6}$ and $[11]_{\bar{3}}$ of the $SU_{c}(3)$ group, while the color configuration of the two-antiquark cluster ($\bar{q}\bar{Q}$) consists of a $[211]_{3}$ triplet and a $[22]_{\bar{6}}$ antisextet.
Thus, a $[222]_{1}$ color singlet for tetraquark states demands the following configurations: $[2]_{6}(q_{1}Q_{2})\otimes[22]_{\bar{6}}(\bar{q}_{3}\bar{Q}_{4})$ and $[11]_{\bar{3}}(q_{1}Q_{2})\otimes[211]_{3}(\bar{q}_{3}\bar{Q}_{4})$, which correspond to the color sextet-antisextet ($6_{c}\otimes\bar{6}_{c}$) and triplet-antitriplet ($\bar{3}_{c}\otimes 3_{c}$) configurations, respectively. The explicit color wave function for each tetraquark color configuration can be found in the previous work Zhao et al. (2021).
$J^{PC}=1^{--}$ Y states could be L=1, S=0 or 2 tetraquark. Thus, two possible spin combinations for tetraquarks are considered as follows: $\left[\psi_{[s=1]}^{qQ}\otimes\psi_{[s=1]}^{\bar{q}\bar{Q}}\right]_{S=0,2}$, and $\psi_{[s=0]}^{qQ}\otimes\psi_{[s=0]}^{\bar{q}\bar{Q}}$.
The complete bases is constructed by coupling among the harmonic oscillator wave functions, may take the general form,
$\psi_{NL}$ $=\sum_{{n_{\chi_{i}},l_{\chi_{i}}}}A(n_{\chi_{1}},n_{\chi_{2}},n_{\chi_{3}},l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}})$ (1)
×$\times\psi_{n_{\chi_{1}}l_{\chi_{1}}}(\vec{\chi}_{1}\,)\otimes\psi_{n_{\chi_{2}}l_{\chi_{2}}}(\vec{\chi}_{2}\,)\otimes\psi_{n_{\chi_{3}}l_{\chi_{3}}}(\vec{\chi}_{3}),$
where $\psi_{n_{\chi_{i}}l_{\chi_{i}}}$ are harmonic oscillator wave functions. The sum ${n_{\chi_{i}},l_{\chi_{i}}}$ is over $n_{\chi_{1}},n_{\chi_{2}},n_{\chi_{3}},l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}$. The relative Jacobi coordinates $\vec{\chi}_{1}$, $\vec{\chi}_{2}$ and $\vec{\chi}_{3}$ are defined as
$\vec{\chi}_{1}=\frac{1}{\sqrt{2}}(\vec{r}_{1}-\vec{r}_{2}),$
$\vec{\chi}_{2}=\frac{1}{\sqrt{2}}(\vec{r}_{3}-\vec{r}_{4}),$
$\vec{\chi}_{3}=\frac{m_{u}\vec{r}_{1}+m_{Q}\vec{r}_{2}}{m_{u}+m_{Q}}-\frac{m_{u}\vec{r}_{3}+m_{Q}\vec{r}_{4}}{m_{u}+m_{Q}},$ (2)
where $\vec{r}_{j}$ are the coordinate of the jth quark. The reduced quark masses are defined as
$u_{1}=\frac{2m_{u}m_{Q}}{m_{u}+m_{Q}},$
$u_{2}=\frac{2m_{u}m_{Q}}{m_{u}+m_{Q}},$
$u_{3}=\frac{(m_{u}+m_{Q})^{2}}{2m_{u}+2m_{Q}},$ (3)
$m_{u}$ and $m_{Q}$ are mass of the light quark and heavy quark respectively. $N$ and $L$ are respectively the total principle quantum number and total angular momentum, and $L=l_{\chi_{1}}+l_{\chi_{2}}+l_{\chi_{3}}=1$ for P wave tetraquark and $N=(2n_{\chi_{1}}+l_{\chi_{1}})+(2n_{\chi_{2}}+l_{\chi_{2}})+(2n_{\chi_{3}}+l_{\chi_{3}})$.
II.2 Hamiltonian
We study meson and tetraquark systems within the non-relativistic quark model Godfrey and Isgur (1985); Schoberl (1986); Barnes et al. (2005). The Hamiltonian has been widely employed in recent years Wang et al. (2021b); Wu and Zhu (2025), taking the form,
$H=$ $H_{0}+H_{so},$
$H_{0}=$ $\sum_{k=1}^{N}(\frac{1}{2}M^{ave}_{k}+\frac{p_{k}^{2}}{2m_{k}})+\sum_{i<j}^{N}(-\frac{3}{16}\vec{\lambda}_{i}\cdot\vec{\lambda}_{j})V_{cen}(r_{ij}),$
$H_{so}=$ $\sum_{i<j}^{N}(-\frac{3}{16}\vec{\lambda}_{i}\cdot\vec{\lambda}_{j})(V_{so}(r_{ij})),$ (4)
where $M^{ave}_{k}$ denotes the spin-averaged mass as $\frac{1}{4}M_{PS}+\frac{3}{4}M_{V}$. $m_{k}$ are the constituent quark masses. $\vec{\lambda}_{i}$ are the quark color operator in SU(3). Cornell-like potential and one-gluon-exchange potential are employed as the central potential, taking the form,
$V_{cen}(r_{ij})=A_{ij}r_{ij}-\frac{B_{ij}}{r_{ij}}+V_{ss}(r_{ij}),$ (5)
where the spin-spin interaction $V_{ss}(r_{ij})$ originates from the one-gluon-exchange potential and takes the form:
$V_{ss}=\frac{1}{6m_{i}m_{j}}\Delta V_{V}(r)\vec{\sigma}_{i}\cdot\vec{\sigma}_{j}=\frac{2B_{ij}\sigma_{ij}^{3}}{3m_{i}m_{j}\sqrt{\pi}}e^{-\sigma_{ij}^{2}r_{ij}^{2}}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j}.$ (6)
The spin-orbital interaction $V_{so}(r_{ij})$ takes the form:
$V_{so}$ $=\frac{1}{r_{ij}}\frac{dV_{V}}{d(r_{ij})}\frac{1}{4}\Bigg{[}\left(\frac{1}{m_{i}^{2}}+\frac{1}{m_{j}^{2}}+\frac{4}{m_{i}m_{j}}\right)\vec{L_{ij}}\cdot\vec{S_{ij}}$
$\quad+\left(\frac{1}{m_{i}^{2}}-\frac{1}{m_{j}^{2}}\right)\vec{L_{ij}}\cdot\left(\vec{s_{i}}-\vec{s_{j}}\right)\Bigg{]}$
$\quad-\frac{1}{r_{ij}}\frac{dV_{S}}{d(r_{ij})}\left(\frac{\vec{L_{ij}}\cdot\vec{s_{i}}}{2m_{i}^{2}}+\frac{\vec{L_{ij}}\cdot\vec{s_{i}}}{m_{j}^{2}}\right)$
$=\left(\frac{1}{m_{i}^{2}}+\frac{1}{m_{j}^{2}}+\frac{4}{m_{i}m_{j}}\right)(\frac{-B_{ij}\sigma_{ij}}{2\sqrt{\pi}})\frac{e^{-\sigma_{ij}^{2}r_{ij}^{2}}}{r_{ij}^{2}}\vec{L_{ij}}\cdot\vec{S}_{ij}$
$+\left(\frac{1}{m_{i}^{2}}+\frac{1}{m_{j}^{2}}+\frac{4}{m_{i}m_{j}}\right)(\frac{-B_{ij}}{4})\frac{Erf[{\sigma_{ij}r_{ij}}]}{r_{ij}^{3}}\vec{L_{ij}}\cdot\vec{S}_{ij}$
$+\left(\frac{-A_{ij}}{2}\right)\frac{1}{r_{ij}}\left(\frac{\vec{L}_{ij}\cdot\vec{s}_{i}}{m_{i}^{2}}+\frac{\vec{L}_{ij}\cdot\vec{s}_{j}}{m_{j}^{2}}\right),$ (7)
where $\vec{\sigma}_{i}$ are the quark spin operator in SU(2). Note that we have employed $V_{V}(r)=-B\,Erf[\sigma r]/r$ and $V_{S}(r)=Ar$, taken from Ref. Schoberl (1986). $m_{i}$ and $m_{j}$ are constituent quark masses of ith and jth quark. $\vec{s_{i}}$ represents the spin operator for ith quark. $\vec{S_{ij}}=\vec{s_{i}}+\vec{s_{j}}$ is the spin operator for (ij)th quark pair. $\vec{L_{ij}}$ is relative orbital operator, taking the form,
$\vec{L_{ij}}=\vec{r_{ij}}\times\vec{p_{ij}}=\vec{L_{ij}}\times\frac{m_{i}\vec{p_{i}}-m_{j}\vec{p_{j}}}{m_{i}+m_{j}}.$ (8)
In line with the previous works, $A_{ij}$, $B_{ij}$, and $\sigma_{ij}$ are proposed to be mass dependent coupling parameters, taking the form
$A_{ij}=a+bm_{ij},\;\;B_{ij}=B_{0}\sqrt{\frac{1}{m_{ij}}},\;\;\sigma_{ij}=\sigma_{0}{m_{ij}}.$ (9)
with $m_{ij}$ being the reduced mass of $i$th and $j$th quarks, defined as $\;m_{ij}=\frac{2m_{i}m_{j}}{m_{i}+m_{j}}$. $a$, $b$, $B_{0}$, and $\sigma_{0}$ are constants. For more detailed discussion, one may refer to Ref. Zhao et al. (2021). The hyperfine coefficient $\sigma_{ij}$ is also proposed to be mass dependent Schoberl (1986).
The central potential $V_{0}$ and the spin-spin interaction $V_{ss}$ are treated as the leading effects, while the remaining spin-orbital interaction $V_{so}$ is considered a perturbation that shifts the mass spectrum. The Schrödinger equation is solved for mesons and tetraquarks within the Hamiltonian $H_{0}$, yielding the eigenvalue $E_{0}$ and eigenstates $\psi_{0}$. The mass spectrum is then calculated by adding the diagonalized $H_{so}$ matrix to the basis of the eigenstates obtained in the previous step, which takes the form,
$E=E_{0}+\langle\psi_{0}|H_{so}|\psi_{0}\rangle.$ (10)
By importing the coupling constants $a$, $b$, and $B_{0}$, as well as the constituent quark masses $m_{u}$, $m_{c}$, and $m_{b}$ from previous work Zhao et al. (2021), the mass spectra of S- and P-wave light, charmed, bottom, charmonium, and bottomonium conventional mesons are calculated in the Hamiltonian in Eq. (II.2). A comparison of the theoretical results, presented in Table 1, with experimental data from the PDG Navas et al. (2024), yields a model coupling parameter $\sigma_{0}=0.85$.
Table 1: S- and P-wave meson states applied to fit the model parameters. The last column shows the deviation between the experimental and theoretical mean values, $D=100\cdot(M^{exp}-M^{cal})/M^{exp}$. $M^{exp}$ taken from PDG Navas et al. (2024).
[…]
$\eta_{b}$ $0^{-+}$ 1S 9399 9383 0.2
2S $9$ 10019 -0.2
$\Upsilon$ $1^{--}$ 1S 9460 9468 -0.1
2S 10023 10055 -0.3
$h_{b}$ $1^{+-}$ 1P 9899 9922 -0.2
2P 10260 10321 -0.6
$\chi_{b0}$ $0^{++}$ 1P 9859 9893 -0.3
2P 10232 10298 -0.6
$\chi_{b1}$ $1^{++}$ 1P 9893 9909 -0.2
2P 10255 10311 -0.5
$\chi_{b2}$ $2^{++}$ 1P 9912 9936 -0.2
2P 10269 10332 -0.6
$\psi$ $1^{--}$ 1S 3097 3097 <0.1
2S 3686 3658 0.7
3S 4040 4030 0.2
$\eta_{c}$ $0^{-+}$ 1S 2984 2927 1.9
2S 3638 3601 1.0
$h_{c}$ $1^{+-}$ 1P 3525 3518 0.2
$\chi_{c0}$ $0^{++}$ 1P 3415 3446 -0.9
2P 3860 3862 <0.1
$\chi_{c1}$ $1^{++}$ 1P 3510 3493 0.5
$\chi_{c2}$ $2^{++}$ 1P 3556 3556 <0.1
2P 3930 3936 -0.1
$B^{0}$ $0^{-}$ 1S 5279 5276 0.1
$B^{*}$ $1^{-}$ 1S 5325 5341 -0.3
$B_{1}$ $1^{+}$ 1P 5721 5800 -1.4
$B_{2}^{*}$ $2^{+}$ 1P 5747 5820 -1.3
$D^{0}$ $0^{-}$ 1S 1865 1870 -0.2
2S 2549 2578 -1.1
$D^{*0}$ $1^{-}$ 1S 2007 2030 -1.2
2S 2627 2626 <0.1
$D_{1}^{0}$ $1^{+}$ 1P 2420 2465 -1.9
$D_{0}^{*0}$ $0^{+}$ 1P 2343 2404 -2.6
$D_{1}^{0}$ $1^{+}$ 1P 2412 2451 -1.6
$D_{2}^{*0}$ $2^{+}$ 1P 2460 2517 -2.3
$\rho$ $1^{--}$ 1S 770 782 -1.5
2S 1450 1401 3.4
$h_{1}$(1170) $1^{+-}$ 1P 1170 1198 -2.4
$b_{1}$(1235) $1^{+-}$ 1P 1235 2.6
$f_{0}$(1370) $0^{++}$ 1P 1200-1500 1137
$f_{1}$(1285) $1^{++}$ 1P 1280 1200 6.2
$a_{1}$(1260) $1^{++}$ 1P 1230 2.4
$f_{2}$(1270) $2^{++}$ 1P 1275 1294 -1.5
$a_{2}$(1320) $2^{++}$ 1P 1318 1.8
II.3 Two-body strong decay
The two-body strong decay properties of the bound states, specifically the Y tetraquark states within the rearrangement mechanism, are studied, as shown in Figure 1. The spin-color factors of the transition amplitudes $T_{cs}$ in the two-body strong decay channels take the form:
$T_{cs}$ $=$ $\langle\psi_{f}^{cs}|\psi_{i}^{cs}\rangle$ (11)
$=$ $\langle\psi_{M_{1}M_{2}}^{c}|\psi_{Y}^{c}\rangle\langle\psi_{M_{1}M_{2}}^{s}|\psi_{Y}^{s}\rangle$
where $\psi_{Y}^{c}$ and $\psi_{Y}^{s}$ are color and spin wave functions of the initial states corresponding to the Y tetraquark states in mass spectrum. $\psi^{c}_{M_{1}M_{2}}$ and $\psi^{s}_{M_{1}M_{2}}$ are the color and spin wave functions of the final states of all possible two-body strong decay channels: $\omega\chi_{c0}$, $\omega\chi_{c1}$, $\omega\chi_{c2}$, $\eta J/\psi$, $\rho J\psi$, $\omega\chi_{b0}$, $\omega\chi_{b1}$, and $\omega\chi_{b2}$.
In the work, the ratio between $|T_{cs}|^{2}$ is considered to be branching ratio approximately supposing that the spatial part from different decay channels provide similar contribution.
Figure 1: Quark rearranged diagram for Y tetraquark in the hidden-charm decay modes.
III Results and Discussion
III.1 Results
The predetermined as well as imported parameters are applied to predict the mass of Y tetraquark states in the Hamiltonian in Eq. (II.2) including the color-spin interaction $V_{ss}$ which may mix up different color-spin configurations.
Because of the cross terms are listed in Appendix B, eigenstates of the Hamiltonian for S=0 are linear combinations of $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(0\otimes 0)}$, $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(1\otimes 1)}$, $\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(0\otimes 0)}$, and $\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(1\otimes 1)}$. Eigenstates for S=2 are linear combinations of $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=2}_{(1\otimes 1)}$ and $\psi^{c}_{6\otimes\bar{6}}\psi^{S=2}_{(1\otimes 1)}$. There is no configuration mixing between S=0 and 2 states, as no cross terms are found. En, where n = 1, 2, 3, $\dots$, represents the $n$th eigenstate of the Hamiltonian in Eq. (II.2) in the notation used later in this work, incorporating color-spin configuration mixing.
The theoretical masses and branching ratios of the E1 to E16 S=0 charmonium-like $Y_{c}$ tetraquarks with various mixed configurations are listed in Table 2. Similarly, the masses and branching ratios of the E1 to E8 mixed S=0 bottomonium-like $Y_{b}$ tetraquarks are listed in Table 3. Likewise, the masses of the E1 to E6 mixed S=2 charmonium-like $Y_{c}$ and bottomonium-like $Y_{b}$ tetraquarks are provided in Table 4.
The spin-color factor of the transition amplitude is zero in the two-body strong decay to $\eta J/\psi$ from S=0 $Y_{c}$ tetraquarks. From S=2 $Y_{c}$ and $Y_{b}$ tetraquarks, spin-color factors in $\omega\chi_{c0}$, $\omega\chi_{c1}$, $\omega\chi_{c2}$, $\eta J/\psi$, $\rho J\psi$, $\omega\chi_{b0}$, $\omega\chi_{b1}$, $\omega\chi_{b2}$ decay channels are all equal to zero.
Table 2: Mass spectrum and branching ratios of mixed S=0 $Y_{c}$ tetraquark states. Mass values are in MeV.
En Mass $\omega\chi_{c0}$ $\omega\chi_{c1}$ $\omega\chi_{c2}$ $\rho J\psi$
E1 4209 1 3 5 9
E2 4258 54 161 269 484
E3 4338 4 12 20 35
E4 4363 36 109 181 326
E5 4369 99 297 496 892
E6 4435 9 27 45 82
E7 4466 19 57 94 170
E8 4480 15 45 75 136
E9 4509 77 230 383 689
E10 4522 17 52 86 155
E11 4534 51 154 257 462
E12 4542 7 20 33 60
E13 4592 131 394 657 1182
E14 4617 66 198 329 593
E15 4650 4 11 19 34
E16 4673 123 368 613 1103
Table 3: Mass spectrum and branching ratios of mixed S=0 $Y_{b}$ tetraquark states. Mass values are in MeV.
En Mass $\omega\chi_{b0}$ $\omega\chi_{b1}$ $\omega\chi_{b2}$
E1 10576 1 3 5
E2 10602 54 161 269
E3 10637 4 12 20
E4 10720 36 109 181
E5 10744 99 297 496
E6 10797 9 27 45
E7 10825 19 57 94
E8 10845 15 45 75
Table 4: Mass spectrum of mixed S=2 $Y_{c}$ tetraquark states. Mass values are in MeV.
En E1 E2 E3 E4 E5 E6
$Y_{c}$ 4400 4431 4446 4560 4580 4615
$Y_{b}$ 10680 10757 10792 10882 10891 10913
Experimental data on $Y_{c}$ and $Y_{b}$ state candidates from different decay processes are listed in Table 5 and will be reviewed and discussed separately below. The assignments of S=0 $Y_{c}$ tetraquarks, along with their theoretical masses and two-body decay branching ratios to $\omega\chi_{c0}$, $\omega\chi_{c1}$, $\omega\chi_{c2}$, and $\rho J/\psi$, are presented in Table 6. Similarly, the assignments of S=0 $Y_{b}$ tetraquarks, along with their theoretical masses and two-body decay branching ratios to $\omega\chi_{b0}$, $\omega\chi_{b1}$, and $\omega\chi_{b2}$, are provided in Table 7.
Table 5: Masses and widths of $Y_{c}$ and $Y_{b}$ state candidates from the cited sources, with units in MeV. PDG represent the names of established states in the PDG, while Assignment refer to the names used in this work. The symbol $\psi$ in the third column represents both $J/\psi$ and $\psi(2S)$.
PDG Assignment $\pi^{+}\pi^{-}\psi$ $K^{+}K^{-}J/\psi$ $\omega\chi_{cJ}$ $\pi^{+}\pi^{-}h_{c}$ $\eta J\psi$ $\pi^{+}\pi^{-}\Upsilon(nS)$ $\omega\chi_{bJ}$
$\psi(4230)$ Y(4230) $4220$, $46$ Ablikim et al. (2020a) $4225$, $73$ Ablikim et al. (2022a) $4230$, $38$ Ablikim et al. (2015) $4218$, $66.0$ Ablikim et al. (2017) $4219$, $82.7$ Ablikim et al. (2020b)
$4221$, $42$ Ablikim et al. (2022b)
$4234$, $18$ Ablikim et al. (2021)
$\psi(4360)$ Y(4340) $4298$, $127$ Ablikim et al. (2022b)
$4340$, $94$ Lees et al. (2014)
$4347$, $103$ Wang et al. (2015)
Y(4380) $4390$, $143$ Ablikim et al. (2021) $4392$, $140$ Ablikim et al. (2017) $4382$, $136$ Ablikim et al. (2020b)
Y(4484) $4484$, $111$ Ablikim et al. (2022a)
Y(4544) $4544$, $116$ Ablikim et al. (2024)
$\psi(4660)$ Y(4660) $4651$, $155$ Ablikim et al. (2021) $4708$, $126$ Ablikim et al. (2023)
$4652$, $68$ Wang et al. (2015)
$4669$, $104$ Lees et al. (2014)
$\Upsilon(10753)$ Y(10753) $10753$, $36$ Mizuk et al. (2019) likely Adachi et al. (2023)
Table 6: Presenting theoretical mass and branching ratio predictions, along with assignments for mixed S=0 $Y_{c}$ tetraquark states. The mass results are given in units of MeV.
En E1 E2 E3 E4 E5 E6 E7 E8 E9 E10 E11 E12 E13 E14 E15 E16
Mass 4209 4258 4338 4363 4369 4435 4466 4480 4509 4522 4534 4542 4592 4617 4650 4673
$\omega\chi_{c0}$ 1 54 4 36 99 9 19 15 77 17 51 7 131 66 4 123
$\omega\chi_{c1}$ 3 161 12 109 297 27 57 45 230 52 154 20 394 198 11 368
$\omega\chi_{c2}$ 5 269 20 181 496 45 94 75 383 86 257 33 657 329 19 613
$\rho J\psi$ 9 484 35 326 892 82 170 136 689 155 462 60 1182 593 34 1103
Data Y(4230) Y’(4230) Y(4340) Y(4380) Y’(4380) Y(4484) Y(4544) Y(4660)
Table 7: Theoretical mass, branching ratio predictions, and assignments for mixed S=0 $Y_{b}$ tetraquark states. Mass values are in MeV.
En E1 E2 E3 E4 E5 E6 E7 E8
Mass 10576 10602 10637 10720 10744 10797 10825 10845
$\omega\chi_{b0}$ 1 54 4 36 99 9 19 15
$\omega\chi_{b1}$ 3 161 12 109 297 27 57 45
$\omega\chi_{b2}$ 5 269 20 181 496 45 94 75
Data Y(10753)
III.2 Y(4230)
In the past two decades, the Y(4230) is an exotic particle observed and confirmed by many experimental collaborations, and is considered part of the broader family of charmonium-like states. First detected in electron-positron annihilation experiments with a mass of approximately 4.23 GeV Aubert et al. (2005), placing it in the region above the open-charm threshold. Unlike conventional quarkonium, which consists of a quark-antiquark pair, the Y(4230) is classified as an exotic hadron due to its unconventional properties, including anomalous decay channels and unexpected production mechanisms. In the previous work, the Y(4230) cannot be accommodated in a conventional charmonium picture including S-D mixing Zhao et al. (2024).
The exact internal structure of the Y(4230) remains a subject of ongoing research, with theoretical models proposing interpretations such as tetraquark states Maiani et al. (2005); Ali et al. (2018); Wang (2019a), molecular bound states of charmed mesonsChiu and Hsieh (2006); Qin et al. (2016), or hybrid mesons Oncala and Soto (2017); Berwein et al. (2015).
Y(4230) has been observed by many decay processes but mainly from electron positron collision, where the processes include $e^{+}e^{-}\to\pi^{+}\pi^{-}\psi$Ablikim et al. (2020a, 2022b, 2021), $\eta J\psi$Ablikim et al. (2020b), $\omega\chi_{c0}$blikim et al. (2015), $\pi^{+}\pi^{-}h_{c}$blikim et al. (2017), and $K^{+}K^{-}J/\psi$Ablikim et al. (2022a), which are listed in Table 5. There are some clues about states identification appearing when the resonances are observed from different channels.
The resonance parameter of Y(4230) in the invariant mass of $\omega\chi_{c0}$blikim et al. (2015) is inconsistent with the line shape of the Y(4260) observed in $\pi^{+}\pi^{-}J/\psi$Aubert et al. (2005) which is the first observed Y states in 2005 by BABAR. This Y(4260) state later superseded by a updated workLees et al. (2012) in PDG because two interfering resonances are not excluded. But both work were from a single-resonance fit.
Since the Y(4230) was also observed in $K^{+}K^{-}J/\psi$Ablikim et al. (2022a) process, and the resonance parameters is consistent with those from Y(4230) observed from $\pi^{+}\pi^{-}J/\psi$ channel, one may classify the Y(4230) from $K^{+}K^{-}J/\psi$Ablikim et al. (2022a) and $\pi^{+}\pi^{-}J/\psi$ to be a same state.
Later, Y(4230) is observed in the process $\pi^{+}\pi^{-}h_{c}$blikim et al. (2017), whose parameters are different from those of Y(4260) observed in $\pi^{+}\pi^{-}J/\psi$ channel in PDG before year 2015. However, the resonance parameters of Y(4230) in $\pi^{+}\pi^{-}h_{c}$blikim et al. (2017) are consistent with those of the resonance observed in $\omega\chi_{c0}$ Ablikim et al. (2015), one may support that the Y(4230) from $\omega\chi_{c0}$ and $\pi^{+}\pi^{-}h_{c}$ are likely a same state.
The discussion above may support to separate Y(4230) into two different states. In assignment Table 6, the one from $\pi^{+}\pi^{-}J/\psi$ and $K^{+}K^{-}J/\psi$ might be tentatively named Y(4230) while the one observed from $\omega\chi_{c0}$ and $\pi^{+}\pi^{-}h_{c}$ might be tentatively named Y’(4230) respectively. The Y(4230) are assigned to be E1 state of P-wave tetraquark with quantum number $1^{--}$ and Y’(4230) are assigned to be E2 state due to their different decay ratio, where E2 state has much larger ratio for $\omega\chi_{c0}$ channel.
III.3 Y(4360)
The Y(4360) is first observed in the process of $e^{+}e^{-}\to\gamma\pi^{+}\pi^{-}\psi(2S)$ Wang et al. (2007). With a mass of approximately 4.36 GeV, Y(4360) resides in the energy region above the open-charm threshold, similar to other Y-states. The Y(4360) is of particular interest due to its unusual decay patterns, including prominent decays into $\psi(2S)$ rather than into lower charmonium states like $J/\psi$, which contrasts with expectations from conventional quarkonium models. This anomalous behavior suggests that Y(4360) could be an exotic hadron.
As a PDG established state, in PDG data listed in Table 5, the mass of $\psi(4360)$ are from 4.3 to 4.4 GeV, which might show the two different states in the large range. One is around 3.4 GeV, and another is around 3.8GeV, which are renamed as Y(4340) and Y(4380) in the assignment respectively.
The Y(4340) observed only in $\pi^{+}\pi^{-}J/\psi$ process is assigned to be E3 tetraquark state taking a small ratio to $\omega\chi_{c0}$ process which is similar to E1 Y(4230) state, one may be the excitation state in same direction with Y(4230) from $\pi^{+}\pi^{-}J/\psi$ process.
Similar with the discussion for Y(4230), Y(4380) is also observed in $\pi^{+}\pi^{-}J/\psi$ Ablikim et al. (2021) and $\pi^{+}\pi^{-}h_{c}$ Ablikim et al. (2017) respectively, where there line shapes is inconsistent with each other Ablikim et al. (2017). One may support that there are two Y states around the mass 4380MeV, which are tentatively named Y(4380) for the one from $\pi^{+}\pi^{-}J/\psi$, and Y’(4380) for the one from $\pi^{+}\pi^{-}h_{c}$.
In the assignment Table 6, E5 state shows similar property with the Y state from $\omega\chi_{c0}$ and $\pi^{+}\pi^{-}h_{c}$, where E5 state has larger ratio to $\omega\chi_{c0}$ process comparing with E4 state, one may assign Y(4380) and Y’(4380) to be E4 and E5 state respectively.
III.4 Y(4660)
The Y(4660) is first discovered in the initial state radiation process $e^{+}e^{-}\to\gamma_{\text{ISR}}\pi^{+}\pi^{-}\psi(2S)$ Lees et al. (2014). With a mass of approximately 4.66 GeV, Y(4660) lies well above the open-charm threshold, adding complexity to the spectrum of charmonium-like states. The Y(4660) exhibits intriguing decay patterns, most notably decaying into $\psi(2S)$ rather than $J/\psi$, a feature shared with other exotic states like Y(4360). This has led to speculation that the Y(4660) could be a tetraquarkLebed and Polosa (2016); Lu et al. (2017); Wang (2019a), a S-D wave mixture stateBhavsar et al. (2018), or a hybrid mesonOncala and Soto (2017); Berwein et al. (2015).
In the assignment Table 6, the Y(4660) is mostly observed in the $\pi^{+}\pi^{-}\psi$ channel, which might be naturally grouped with Y(4230), Y(4340) and Y(4380) from the same process. One may assign it to be E15 tetraquark state which has a small decay ratio to $\omega\chi_{c0}$.
III.5 Other $Y_{c}$ states
[…]
The theoretical results of branching ratio for $1^{--}$ Y tetraquark state decaying to $\eta J/\psi$ are all equal to zero. Thus, the Y(4230) state and the Y(4360) state observed from $\eta J/\psi$ Ablikim et al. (2020b) can not be accommodated in the current tetraquark picture, and these two states could be other structure, for example, hadronic molecule states Chiu and Hsieh (2006); Qin et al. (2016); Ding (2009); Close and Downum (2009); Close et al. (2010), or hybrid mesons Liu et al. (2012); Luo and Liu (2006); Zhu (2005); Oncala and Soto (2017); Berwein et al. (2015).
III.6 Y(10753)
The Y(10753) bottomonium-like state was firstly observed by Belle in the $e^{+}e^{-}\to\pi^{+}\pi^{-}\Upsilon(nS)$ process Mizuk et al. (2019). Recently, the Belle II Collaboration reported another observation that strongly supports the existence of the radiative transition process $\Upsilon(10753)\to\omega\chi_{bJ}(1P)$ Adachi et al. (2023). The energy dependence of the Born cross sections for Y(10753) in the $\omega\chi_{bJ}(1P)$ channel Adachi et al. (2023) was found to be consistent with the shape of the Y(10753) state in the $\pi^{+}\pi^{-}\Upsilon(nS)$ channel Mizuk et al. (2019). Belle II concluded that the internal structure of Y(10753) may differ from that of $\Upsilon(10860)$, where the latter is well understood as a predominantly conventional bottomonium state.
In previous work, the $Y(10753)$ could not be accommodated within a conventional bottomonium picture, including S-D mixing Zhao et al. (2024). After the experimental discovery, a P-wave hidden-bottom tetraquark interpretation was proposed in the first set of theoretical pictures Wang (2019b). For a tetraquark mixing interpretation, one may refer to Ref Ali et al. (2020).
In the assignment Table 7, the Y(10753) observed in $\pi^{+}\pi^{-}\Upsilon(nS)$ and $\omega\chi_{bJ}(1P)$ channel is assigned to be E5 tetraquark state due to mass agreement and a large ratio to $\omega\chi_{bJ}$.
IV Summary
The masses of P-wave charmonium-like tetraquark states are calculated using a constituent quark model (CQM) that incorporates a Cornell-like potential and the Breit-Fermi interaction. All model parameters were predetermined by reproducing the mass spectra of low-lying S- and P-wave light, charmed, and charmonium mesons. The theoretical results for P-wave tetraquarks are compared with selected exotic states, known as Y states, and a tentative assignment is proposed.
The PDG established state around 4.23 GeV may split into two distinct states. In the assignment, the Y(4230), observed in the $\pi^{+}\pi^{-}J/\psi$ and $K^{+}K^{-}J/\psi$ channels, is assigned as E1 state of an S=0 $1^{--}$ P-wave tetraquark. The Y’(4230), observed in the $\omega\chi_{c0}$ and $\pi^{+}\pi^{-}h_{c}$ channels, is assigned as E2 state of an S=0 $1^{--}$ P-wave tetraquark.
The PDG established state around 4.36 GeV may also split into three distinct states. In the assignment, the Y(4340) observed only in the $\pi^{+}\pi^{-}J/\psi$ process, is assigned as E3 state of an S=0 $1^{--}$ tetraquark, which can be paired with the E1 Y(4230) state. The Y(4380), observed in the $\pi^{+}\pi^{-}J/\psi$ channel, is assigned as E4 state of an S=0 $1^{--}$ tetraquark, while the Y’(4380), observed in the $\pi^{+}\pi^{-}h_{c}$ channel, is assigned as E5 state of an S=0 $1^{--}$ tetraquark.
The PDG established state Y(4660), observed in the $\pi^{+}\pi^{-}J/\psi$ channel and grouped with Y(4230), Y(4340) and Y(4380) from the same process, is assigned as the E15 S=0 $1^{--}$ tetraquark state.
The Y(4484), observed in the $K^{+}K^{-}J/\psi$ channel by BESIII, is grouped with Y(4230) from the $\pi^{+}\pi^{-}J/\psi$ channel and tentatively assigned as the E8 S=0 $1^{--}$ tetraquark state.
The Y(4544), newly observed in the $\omega\chi_{c1}$ channel by BESIII, is grouped with Y’(4230) from the $\omega\chi_{c0}$ channel and tentatively assigned as E11 S=0 $1^{--}$ tetraquark state.
The Y(10753), observed in $\pi^{+}\pi^{-}\Upsilon(nS)$ and $\omega\chi_{bJ}(1P)$ channel by Belle, is assigned as an E5 S=0 $1^{--}$ bottomonium-like tetraquark state.
Acknowledgements.
This work was supported by (i) Suranaree University of Technology (SUT), (ii) Thailand Science Research and Innovation (TSRI), and (iii) National Science Research and Innovation Fund (NSRF), Project No. 179349.
Appendix A Tetraquark spatial wave function
The total spatial wave function of tetraquark, coupling among the $\chi_{1}$, $\chi_{2}$ and $\chi_{3}$ harmonic oscillator wave functions, may take the general form,
$\psi_{NL}$ $=\sum_{{n_{\chi_{i}},l_{\chi_{i}}}}A(n_{\chi_{1}},n_{\chi_{2}},n_{\chi_{3}},l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}})$ (12)
×$\times\psi_{n_{\chi_{1}}l_{\chi_{1}}}(\vec{\chi}_{1}\,)\otimes\psi_{n_{\chi_{2}}l_{\chi_{2}}}(\vec{\chi}_{2}\,)\otimes\psi_{n_{\chi_{3}}l_{\chi_{3}}}(\vec{\chi}_{3})$
where $\psi_{n_{\chi_{i}}l_{\chi_{i}}}$ are harmonic oscillator wave functions. The sum ${n_{\chi_{i}},l_{\chi_{i}}}$ is over $n_{\chi_{1}},n_{\chi_{2}},n_{\chi_{3}},l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}$. $N$ and $L$ are respectively the total principle quantum number and total angular momentum, and $L=l_{\chi_{1}}+l_{\chi_{2}}+l_{\chi_{3}}=1$ for P wave tetraquark and $N=(2n_{\chi_{1}}+l_{\chi_{1}})+(2n_{\chi_{2}}+l_{\chi_{2}})+(2n_{\chi_{3}}+l_{\chi_{3}})$.
The complete bases of the tetraquarks are listed in Table 8 up to $N=3$ for reference.
Table 8: The complete bases of tetraquark with quantum number, $N=2n+L$ and $L=1$.
$\psi_{11}$ $\psi_{0,0}(\vec{\chi}_{1}\,)\psi_{0,0}(\vec{\chi}_{2}\,)\psi_{0,1}(\vec{\chi}_{3}\,)$, $\psi_{0,0}(\vec{\chi}_{1}\,)\psi_{0,1}(\vec{\chi}_{2}\,)\psi_{0,0}(\vec{\chi}_{3}\,)$,
$\psi_{0,1}(\vec{\chi}_{1}\,)\psi_{0,0}(\vec{\chi}_{2}\,)\psi_{0,0}(\vec{\chi}_{3}\,)$
$\psi_{31}$ $\psi_{1,0}(\vec{\chi}_{1}\,)\psi_{0,0}(\vec{\chi}_{2}\,)\psi_{0,1}(\vec{\chi}_{3}\,)$, $\psi_{1,0}(\vec{\chi}_{1}\,)\psi_{0,1}(\vec{\chi}_{2}\,)\psi_{0,0}(\vec{\chi}_{3}\,)$,
$\psi_{1,1}(\vec{\chi}_{1}\,)\psi_{0,0}(\vec{\chi}_{2}\,)\psi_{0,0}(\vec{\chi}_{3}\,)$, $\psi_{0,0}(\vec{\chi}_{1}\,)\psi_{1,0}(\vec{\chi}_{2}\,)\psi_{0,1}(\vec{\chi}_{3}\,)$,
$\psi_{0,0}(\vec{\chi}_{1}\,)\psi_{1,1}(\vec{\chi}_{2}\,)\psi_{0,0}(\vec{\chi}_{3}\,)$, $\psi_{0,1}(\vec{\chi}_{1}\,)\psi_{1,0}(\vec{\chi}_{2}\,)\psi_{0,0}(\vec{\chi}_{3}\,)$,
$\psi_{0,0}(\vec{\chi}_{1}\,)\psi_{0,0}(\vec{\chi}_{2}\,)\psi_{1,1}(\vec{\chi}_{3}\,)$, $\psi_{0,0}(\vec{\chi}_{1}\,)\psi_{0,1}(\vec{\chi}_{2}\,)\psi_{1,0}(\vec{\chi}_{3}\,)$,
$\psi_{0,1}(\vec{\chi}_{1}\,)\psi_{0,0}(\vec{\chi}_{2}\,)\psi_{1,0}(\vec{\chi}_{3}\,)$
Appendix B Cross term between different configuration from spin-spin interaction
Table 9: Expectation values of the color-spin operator for the S=0 tetraquark state. $<\psi^{k1}_{c}\psi^{h1}_{s}|\vec{\lambda}_{i}\vec{\lambda}_{j}\vec{\sigma}_{i}\vec{\sigma}_{j}|\psi^{k2}_{c}\psi^{h2}_{s}>$
$\vec{\lambda}_{i}\vec{\lambda}_{j}\vec{\sigma}_{i}\vec{\sigma}_{j}$ $|\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(0\otimes 0)}\rangle$ $|\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(1\otimes 1)}\rangle$ $|\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(0\otimes 0)}\rangle$ $|\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(1\otimes 1)}\rangle$
$|\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(0\otimes 0)}\rangle$ $\left(8,0,0,0,0,8\right)$ $\left(0,\frac{4}{\sqrt{3}},-\frac{4}{\sqrt{3}},-\frac{4}{\sqrt{3}},\frac{4}{\sqrt{3}},0\right)$ $\left(0,0,0,0,0,0\right)$ $\left(0,2\sqrt{6},2\sqrt{6},2\sqrt{6},2\sqrt{6},0\right)$
$|\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(1\otimes 1)}\rangle$ $\left(0,\frac{4}{\sqrt{3}},-\frac{4}{\sqrt{3}},-\frac{4}{\sqrt{3}},\frac{4}{\sqrt{3}},0\right)$ $\left(-\frac{8}{3},\frac{8}{3},\frac{8}{3},\frac{8}{3},\frac{8}{3},-\frac{8}{3}\right)$ $\left(0,2\sqrt{6},2\sqrt{6},2\sqrt{6},2\sqrt{6},0\right)$ $\left(0,4\sqrt{2},-4\sqrt{2},-4\sqrt{2},4\sqrt{2},0\right)$
$|\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(0\otimes 0)}\rangle$ $\left(0,0,0,0,0,0\right)$ $\left(0,2\sqrt{6},2\sqrt{6},2\sqrt{6},2\sqrt{6},0\right)$ $\left(-4,0,0,0,0,-4\right)$ $\left(0,\frac{10}{\sqrt{3}},-\frac{10}{\sqrt{3}},-\frac{10}{\sqrt{3}},\frac{10}{\sqrt{3}},0\right)$
$|\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(1\otimes 1)}\rangle$ $\left(0,2\sqrt{6},2\sqrt{6},2\sqrt{6},2\sqrt{6},0\right)$ $\left(0,4\sqrt{2},-4\sqrt{2},-4\sqrt{2},4\sqrt{2},0\right)$ $\left(0,\frac{10}{\sqrt{3}},-\frac{10}{\sqrt{3}},-\frac{10}{\sqrt{3}},\frac{10}{\sqrt{3}},0\right)$ $\left(\frac{4}{3},\frac{20}{3},\frac{20}{3},\frac{20}{3},\frac{20}{3},\frac{4}{3}\right)$
Table 10: Expectation values of the color-spin operator for S=2 tetraquark state. $<\psi^{k1}_{c}\psi^{h1}_{s}|\vec{\lambda}_{i}\vec{\lambda}_{j}\vec{\sigma}_{i}\vec{\sigma}_{j}|\psi^{k2}_{c}\psi^{h2}_{s}>$
$\vec{\lambda}_{i}\vec{\lambda}_{j}\vec{\sigma}_{i}\vec{\sigma}_{j}$ $|\psi^{c}_{\bar{3}\otimes 3}\psi^{S=2}_{(1\otimes 1)}\rangle$ $|\psi^{c}_{6\otimes\bar{6}}\psi^{S=2}_{(1\otimes 1)}\rangle$
$|\psi^{c}_{\bar{3}\otimes 3}\psi^{S=2}_{(1\otimes 1)}\rangle$ $\left(-\frac{8}{3},\frac{4}{3},\frac{4}{3},\frac{4}{3},\frac{4}{3},-\frac{8}{3}\right)$ $\left(0,4\sqrt{2},-4\sqrt{2},-4\sqrt{2},4\sqrt{2},0\right)$
$|\psi^{c}_{6\otimes\bar{6}}\psi^{S=2}_{(1\otimes 1)}\rangle$ $\left(0,4\sqrt{2},-4\sqrt{2},-4\sqrt{2},4\sqrt{2},0\right)$ $\left(\frac{4}{3},-\frac{10}{3},-\frac{10}{3},-\frac{10}{3},-\frac{10}{3},\frac{4}{3}\right)$
The color-spin operator $\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j}$ in the Hamiltonian in Eq. (II.2), along with the mass-dependent factor in the spin-spin interaction,
$\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}V_{ss}$ $=\frac{2B_{ij}\sigma_{ij}^{3}}{3m_{i}m_{j}\sqrt{\pi}}e^{-\sigma_{ij}^{2}r_{ij}^{2}}\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j},$ (13)
mixes different color-spin configurations, leading to nonzero cross terms for charmonium-like and bottomonium-like tetraquarks.
For reference, the expectation values of $O_{ij}=\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j}$ for S=0 and S=2 states are listed in Table 9 and Table 10 respectively, with all components given in the format $(O_{12},O_{13},O_{23},O_{14},O_{24},O_{34})$. The mass-dependent factor in Eq. (13) induces nonzero cross terms between different color-spin configurations.
In conclusion, the S=0 charmonium-like and bottomonium-like tetraquarks are linear combinations of $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(0\otimes 0)}$, $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(1\otimes 1)}$, $\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(0\otimes 0)}$, and $\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(1\otimes 1)}$. Similarly, the S=2 charmonium-like and bottomonium-like tetraquarks are linear combinations of $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=2}_{(1\otimes 1)}$ and $\psi^{c}_{6\otimes\bar{6}}\psi^{S=2}_{(1\otimes 1)}$. There is no configuration mixing between S=0 and 2 states, as no cross terms are found.

### v3

@@
 All model parameters were imported from previous work, and predetermined by studying the low-lying conventional S- and P-wave light, charmed, bottom, charmonium, and bottomonium meson mass spectra.
+The lowest $1^{--}$ tetraquark mass is predicted to be around 4.15 GeV.
 The decay widths of $1^{--}$ P-wave tetraquark states are calculated for possible two-body strong decay channels within the rearrangement mechanism, including $\omega\chi_{cJ}$, $\eta J/\psi$, and $\rho J/\psi$ for charmonium-like tetraquarks, and $\omega\chi_{bJ}$ for bottomonium-like tetraquarks.
 The tetraquark theoretical results are compared with the selected exotic states, also known as Y states, and tentative assignments are suggested.
-This study suggests that $\psi(4230)$, $\psi(4360)$, $\psi(4660)$, and $\Upsilon$(10753) may be P-wave tetraquark states and that multiple states might exist around 4.23 GeV and 4.36 GeV.
+This study suggests that $\psi(4230)$, $\psi(4360)$, $\psi(4660)$, and $\Upsilon$(10753) may be P-wave tetraquark states and that multiple states might exist around 4.36 GeV.
 I Introduction
@@
 II.1 Quark configurations of tetraquark
-The construction of tetraquark states follows to the principle that a tetraquark state must be a color singlet, meaning the color wave function of the tetraquark must be a $[222]_{1}$ singlet within the $SU_{c}(3)$ group.
+The construction of tetraquark states follows to the principle that a tetraquark state must be a color singlet, meaning the color wave function of the tetraquark must be a $222]_{1}$ singlet within the $SU_{c}(3)$ group.
 In this work we consider only charmonium-like and bottomonium-like ($q_{1}Q_{2}\bar{q}_{3}\bar{Q}_{4}$) states, where $q_{1}$ and $Q_{2}$ are light quark and heavy quark, and $\bar{q}_{3}$ and $\bar{Q}_{4}$ are light antiquark and heavy antiquark respectively.
-The permutation symmetry of the two-quark cluster ($qQ$) in tetraquark states is described by the Young tabloids $[2]_{6}$ and $[11]_{\bar{3}}$ of the $SU_{c}(3)$ group, while the color configuration of the two-antiquark cluster ($\bar{q}\bar{Q}$) consists of a $[211]_{3}$ triplet and a $[22]_{\bar{6}}$ antisextet.
-Thus, a $[222]_{1}$ color singlet for tetraquark states demands the following configurations: $[2]_{6}(q_{1}Q_{2})\otimes[22]_{\bar{6}}(\bar{q}_{3}\bar{Q}_{4})$ and $[11]_{\bar{3}}(q_{1}Q_{2})\otimes[211]_{3}(\bar{q}_{3}\bar{Q}_{4})$, which correspond to the color sextet-antisextet ($6_{c}\otimes\bar{6}_{c}$) and triplet-antitriplet ($\bar{3}_{c}\otimes 3_{c}$) configurations, respectively.
+The permutation symmetry of the two-quark cluster ($qQ$) in tetraquark states is described by the Young tabloids $2]_{6}$ and $11]_{\bar{3}}$ of the $SU_{c}(3)$ group, while the color configuration of the two-antiquark cluster ($\bar{q}\bar{Q}$) consists of a $211]_{3}$ triplet and a $22]_{\bar{6}}$ antisextet.
+Thus, a $222]_{1}$ color singlet for tetraquark states demands the following configurations: $2]_{6}(q_{1}Q_{2})\otimes[22]_{\bar{6}}(\bar{q}_{3}\bar{Q}_{4})$ and $11]_{\bar{3}}(q_{1}Q_{2})\otimes[211]_{3}(\bar{q}_{3}\bar{Q}_{4})$, which correspond to the color sextet-antisextet ($6_{c}\otimes\bar{6}_{c}$) and triplet-antitriplet ($\bar{3}_{c}\otimes 3_{c}$) configurations, respectively.
 The explicit color wave function for each tetraquark color configuration can be found in the previous work Zhao et al.
@@
 $\psi_{NL}$ $=\sum_{{n_{\chi_{i}},l_{\chi_{i}}}}A(n_{\chi_{1}},n_{\chi_{2}},n_{\chi_{3}},l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}})$ (1)
-×$\times\psi_{n_{\chi_{1}}l_{\chi_{1}}}(\vec{\chi}_{1}\,)\otimes\psi_{n_{\chi_{2}}l_{\chi_{2}}}(\vec{\chi}_{2}\,)\otimes\psi_{n_{\chi_{3}}l_{\chi_{3}}}(\vec{\chi}_{3}),$
+$\times\psi_{n_{\chi_{1}}l_{\chi_{1}}}(\vec{\chi}_{1}\,)\otimes\psi_{n_{\chi_{2}}l_{\chi_{2}}}(\vec{\chi}_{2}\,)\otimes\psi_{n_{\chi_{3}}l_{\chi_{3}}}(\vec{\chi}_{3}),$
 where $\psi_{n_{\chi_{i}}l_{\chi_{i}}}$ are harmonic oscillator wave functions.
@@
 $u_{2}=\frac{2m_{u}m_{Q}}{m_{u}+m_{Q}},$
-$u_{3}=\frac{(m_{u}+m_{Q})^{2}}{2m_{u}+2m_{Q}},$ (3)
+$u_{3}=\frac{m_{u}+m_{Q}}{2},$ (3)
 $m_{u}$ and $m_{Q}$ are mass of the light quark and heavy quark respectively.
@@
 $\vec{L_{ij}}$ is relative orbital operator, taking the form,
-$\vec{L_{ij}}=\vec{r_{ij}}\times\vec{p_{ij}}=\vec{L_{ij}}\times\frac{m_{i}\vec{p_{i}}-m_{j}\vec{p_{j}}}{m_{i}+m_{j}}.$ (8)
+$\vec{L_{ij}}=\vec{r_{ij}}\times\vec{p_{ij}}=\vec{r_{ij}}\times\frac{m_{i}\vec{p_{i}}-m_{j}\vec{p_{j}}}{m_{i}+m_{j}}.$ (8)
 In line with the previous works, $A_{ij}$, $B_{ij}$, and $\sigma_{ij}$ are proposed to be mass dependent coupling parameters, taking the form
@@
 A comparison of the theoretical results, presented in Table 1, with experimental data from the PDG Navas et al.
-(2024), yields a model coupling parameter $\sigma_{0}=0.85$.
+(2024), yields a model coupling parameter $\sigma_{0}=0.7$.
 Table 1: S- and P-wave meson states applied to fit the model parameters.
@@
 (2024).
-[…]
-$\eta_{b}$ $0^{-+}$ 1S 9399 9383 0.2
-2S $9$ 10019 -0.2
-$\Upsilon$ $1^{--}$ 1S 9460 9468 -0.1
-2S 10023 10055 -0.3
+Meson $J^{P(C)}$ nL $M^{exp}{\rm(MeV)}$ $M^{cal}{\rm(MeV)}$ D ($\%$)
+$\eta_{b}$ $0^{-+}$ 1S 9399 9394 0.1
+2S $99$ 10024 -0.3
+$\Upsilon$ $1^{--}$ 1S 9460 9467 -0.1
+2S 10023 10054 -0.3
 $h_{b}$ $1^{+-}$ 1P 9899 9922 -0.2
-2P 10260 10321 -0.6
-$\chi_{b0}$ $0^{++}$ 1P 9859 9893 -0.3
-2P 10232 10298 -0.6
-$\chi_{b1}$ $1^{++}$ 1P 9893 9909 -0.2
-2P 10255 10311 -0.5
-$\chi_{b2}$ $2^{++}$ 1P 9912 9936 -0.2
+2P 10260 10320 -0.6
+$\chi_{b0}$ $0^{++}$ 1P 9859 9896 -0.4
+2P 10232 10301 -0.7
+$\chi_{b1}$ $1^{++}$ 1P 9893 9911 -0.2
+2P 10255 10312 -0.6
+$\chi_{b2}$ $2^{++}$ 1P 9912 9935 -0.2
 2P 10269 10332 -0.6
-$\psi$ $1^{--}$ 1S 3097 3097 <0.1
-2S 3686 3658 0.7
-3S 4040 4030 0.2
-$\eta_{c}$ $0^{-+}$ 1S 2984 2927 1.9
-2S 3638 3601 1.0
-$h_{c}$ $1^{+-}$ 1P 3525 3518 0.2
-$\chi_{c0}$ $0^{++}$ 1P 3415 3446 -0.9
-2P 3860 3862 <0.1
-$\chi_{c1}$ $1^{++}$ 1P 3510 3493 0.5
-$\chi_{c2}$ $2^{++}$ 1P 3556 3556 <0.1
-2P 3930 3936 -0.1
-$B^{0}$ $0^{-}$ 1S 5279 5276 0.1
-$B^{*}$ $1^{-}$ 1S 5325 5341 -0.3
+$\eta_{c}$ $0^{-+}$ 1S 2984 2962 0.7
+2S 3638 3614 0.7
+$\psi$ $1^{--}$ 1S 3097 3090 0.2
+2S 3686 3655 0.8
+3S $40$ 4027 0.3
+$h_{c}$ $1^{+-}$ 1P 3525 3516 0.2
+$\chi_{c0}$ $0^{++}$ 1P 3415 3464 -1.4
+2P 3860 3876 -0.4
+$\chi_{c1}$ $1^{++}$ 1P 3510 3500 0.3
+$\chi_{c2}$ $2^{++}$ 1P 3556 3553 0.1
+2P 3930 3932 -0.1
+$B^{0}$ $0^{-}$ 1S 5279 5289 -0.2
+$B^{*}$ $1^{-}$ 1S 5325 5337 -0.2
 $B_{1}$ $1^{+}$ 1P 5721 5800 -1.4
-$B_{2}^{*}$ $2^{+}$ 1P 5747 5820 -1.3
-$D^{0}$ $0^{-}$ 1S 1865 1870 -0.2
-2S 2549 2578 -1.1
-$D^{*0}$ $1^{-}$ 1S 2007 2030 -1.2
-2S 2627 2626 <0.1
+$B_{2}^{*}$ $2^{+}$ 1P 5747 5814 -1.2
+$D^{0}$ $0^{-}$ 1S 1865 1906 -2.2
+2S 2549 2590 -1.6
+$D^{*0}$ $1^{-}$ 1S 2007 2021 -0.7
+2S 2627 2622 0.2
 $D_{1}^{0}$ $1^{+}$ 1P 2420 2465 -1.9
-$D_{0}^{*0}$ $0^{+}$ 1P 2343 2404 -2.6
-$D_{1}^{0}$ $1^{+}$ 1P 2412 2451 -1.6
-$D_{2}^{*0}$ $2^{+}$ 1P 2460 2517 -2.3
-$\rho$ $1^{--}$ 1S 770 782 -1.5
-2S 1450 1401 3.4
-$h_{1}$(1170) $1^{+-}$ 1P 1170 1198 -2.4
-$b_{1}$(1235) $1^{+-}$ 1P 1235 2.6
-$f_{0}$(1370) $0^{++}$ 1P 1200-1500 1137
-$f_{1}$(1285) $1^{++}$ 1P 1280 1200 6.2
-$a_{1}$(1260) $1^{++}$ 1P 1230 2.4
-$f_{2}$(1270) $2^{++}$ 1P 1275 1294 -1.5
-$a_{2}$(1320) $2^{++}$ 1P 1318 1.8
+$D_{0}^{*0}$ $0^{+}$ 1P 2343 2438 -4.1
+$D_{1}^{0}$ $1^{+}$ 1P 2412 2465 -2.2
+$D_{2}^{*0}$ $2^{+}$ 1P 2460 2508 -2.0
+$\rho$ $1^{--}$ 1S 770 766 0.5
+2S 1450 1395 3.8
+$h_{1}$(1170) $1^{+-}$ 1P 1170 1202 -2.7
+$b_{1}$(1235) $1^{+-}$ 1P 1235 2.3
+$f_{0}$(1370) $0^{++}$ 1P 1200-1500 1191
+$f_{1}$(1285) $1^{++}$ 1P 1280 1223 4.4
+$a_{1}$(1260) $1^{++}$ 1P 1230 0.6
+$f_{2}$(1270) $2^{++}$ 1P 1275 1277 -0.2
+$a_{2}$(1320) $2^{++}$ 1P 1318 3.1
 II.3 Two-body strong decay
 The two-body strong decay properties of the bound states, specifically the Y tetraquark states within the rearrangement mechanism, are studied, as shown in Figure 1.
-The spin-color factors of the transition amplitudes $T_{cs}$ in the two-body strong decay channels take the form:
 $T_{cs}$ $=$ $\langle\psi_{f}^{cs}|\psi_{i}^{cs}\rangle$ (11)
@@
 The predetermined as well as imported parameters are applied to predict the mass of Y tetraquark states in the Hamiltonian in Eq.
-(II.2) including the color-spin interaction $V_{ss}$ which may mix up different color-spin configurations.
-Because of the cross terms are listed in Appendix B, eigenstates of the Hamiltonian for S=0 are linear combinations of $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(0\otimes 0)}$, $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(1\otimes 1)}$, $\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(0\otimes 0)}$, and $\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(1\otimes 1)}$.
+(II.2) including the color dependent central potential $V_{cen}$ which may mix up different color-spin configurations.
+Because of the cross terms from color operator $(\vec{\lambda}_{i}\cdot\vec{\lambda}_{j})$ and spin operator $(\vec{\sigma}_{i}\cdot\vec{\sigma}_{j})$, listed in Appendix B, eigenstates of the Hamiltonian for S=0 are linear combinations of $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(0\otimes 0)}$, $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(1\otimes 1)}$, $\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(0\otimes 0)}$, and $\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(1\otimes 1)}$.
 Eigenstates for S=2 are linear combinations of $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=2}_{(1\otimes 1)}$ and $\psi^{c}_{6\otimes\bar{6}}\psi^{S=2}_{(1\otimes 1)}$.
@@
 En Mass $\omega\chi_{c0}$ $\omega\chi_{c1}$ $\omega\chi_{c2}$ $\rho J\psi$
-E1 4209 1 3 5 9
-E2 4258 54 161 269 484
-E3 4338 4 12 20 35
-E4 4363 36 109 181 326
-E5 4369 99 297 496 892
-E6 4435 9 27 45 82
-E7 4466 19 57 94 170
-E8 4480 15 45 75 136
-E9 4509 77 230 383 689
-E10 4522 17 52 86 155
-E11 4534 51 154 257 462
-E12 4542 7 20 33 60
+E1 4153 1 3 5 9
+E2 4214 54 161 269 484
+E3 4313 4 12 20 35
+E4 4315 36 109 181 326
+E5 4391 99 297 496 892
+E6 4413 9 27 45 82
+E7 4488 19 57 94 170
+E8 4496 15 45 75 136
+E9 4518 77 230 383 689
+E10 4536 17 52 86 155
+E11 4562 51 154 257 462
+E12 4584 7 20 33 60
 E13 4592 131 394 657 1182
-E14 4617 66 198 329 593
-E15 4650 4 11 19 34
-E16 4673 123 368 613 1103
+E14 4604 66 198 329 593
+E15 4621 4 11 19 34
+E16 4662 123 368 613 1103
 Table 3: Mass spectrum and branching ratios of mixed S=0 $Y_{b}$ tetraquark states. Mass values are in MeV.
 En Mass $\omega\chi_{b0}$ $\omega\chi_{b1}$ $\omega\chi_{b2}$
-E1 10576 1 3 5
-E2 10602 54 161 269
-E3 10637 4 12 20
-E4 10720 36 109 181
-E5 10744 99 297 496
-E6 10797 9 27 45
-E7 10825 19 57 94
-E8 10845 15 45 75
+E1 10497 1 3 5
+E2 10542 54 161 269
+E3 10573 4 12 20
+E4 10685 36 109 181
+E5 10749 99 297 496
+E6 10769 9 27 45
+E7 10790 19 57 94
+E8 10808 15 45 75
 Table 4: Mass spectrum of mixed S=2 $Y_{c}$ tetraquark states. Mass values are in MeV.
 En E1 E2 E3 E4 E5 E6
-$Y_{c}$ 4400 4431 4446 4560 4580 4615
-$Y_{b}$ 10680 10757 10792 10882 10891 10913
+$Y_{c}$ 4300 4347 4467 4574 4575 4614
+$Y_{b}$ 10568 10665 10791 10828 10851 10941
 Experimental data on $Y_{c}$ and $Y_{b}$ state candidates from different decay processes are listed in Table 5 and will be reviewed and discussed separately below.
-The assignments of S=0 $Y_{c}$ tetraquarks, along with their theoretical masses and two-body decay branching ratios to $\omega\chi_{c0}$, $\omega\chi_{c1}$, $\omega\chi_{c2}$, and $\rho J/\psi$, are presented in Table 6.
+The assignments of S=0 and S=2 $Y_{c}$ tetraquarks, along with their theoretical masses and two-body decay branching ratios to $\omega\chi_{c0}$, $\omega\chi_{c1}$, $\omega\chi_{c2}$, and $\rho J/\psi$, are presented in Table 6.
 Similarly, the assignments of S=0 $Y_{b}$ tetraquarks, along with their theoretical masses and two-body decay branching ratios to $\omega\chi_{b0}$, $\omega\chi_{b1}$, and $\omega\chi_{b2}$, are provided in Table 7.
@@
 (2020a) $4225$, $73$ Ablikim et al.
-(2022a) $4230$, $38$ Ablikim et al.
+(2022a) $4218$, $28$ Ablikim et al.
 (2019) $4218$, $66.0$ Ablikim et al.
-(2017) $4219$, $82.7$ Ablikim et al.
+(2017a) $4219$, $82.7$ Ablikim et al.
 (2020b)
@@
 $\psi(4360)$ Y(4300) $4298$, $127$ Ablikim et al. (2022b)
-$4340$, $94$ Lees et al. (2014)
+Y(4340) $4340$, $94$ Lees et al. (2014)
 $4347$, $103$ Wang et al. (2015)
-Y(4380) $4390$, $143$ Ablikim et al. (2021) $4392$, $140$ Ablikim et al. (2017) $4382$, $136$ Ablikim et al. (2020b)
+Y(4390) $4390$, $143$ Ablikim et al. (2021) $4392$, $140$ Ablikim et al. (2017a) $4382$, $136$ Ablikim et al. (2020b)
 Y(4484) $4484$, $111$ Ablikim et al. (2022a)
@@
 $4669$, $104$ Lees et al. (2014)
-$\Upsilon(10753)$ Y(10753) $10753$, $36$ Mizuk et al. (2019) likely Adachi et al. (2023)
-Table 6: Presenting theoretical mass and branching ratio predictions, along with assignments for mixed S=0 $Y_{c}$ tetraquark states. The mass results are given in units of MeV.
-En E1 E2 E3 E4 E5 E6 E7 E8 E9 E10 E11 E12 E13 E14 E15 E16
-Mass 4209 4258 4338 4363 4369 4435 4466 4480 4509 4522 4534 4542 4592 4617 4650 4673
-$\omega\chi_{c0}$ 1 54 4 36 99 9 19 15 77 17 51 7 131 66 4 123
-$\omega\chi_{c1}$ 3 161 12 109 297 27 57 45 230 52 154 20 394 198 11 368
-$\omega\chi_{c2}$ 5 269 20 181 496 45 94 75 383 86 257 33 657 329 19 613
-$\rho J\psi$ 9 484 35 326 892 82 170 136 689 155 462 60 1182 593 34 1103
-Data Y(4230) Y’(4230) Y(4340) Y(4380) Y’(4380) Y(4484) Y(4544) Y(4660)
-Table 7: Theoretical mass, branching ratio predictions, and assignments for mixed S=0 $Y_{b}$ tetraquark states. Mass values are in MeV.
-En E1 E2 E3 E4 E5 E6 E7 E8
-Mass 10576 10602 10637 10720 10744 10797 10825 10845
-$\omega\chi_{b0}$ 1 54 4 36 99 9 19 15
-$\omega\chi_{b1}$ 3 161 12 109 297 27 57 45
-$\omega\chi_{b2}$ 5 269 20 181 496 45 94 75
+$\Upsilon(10753)$ Y(10753) $10753$, $36$ Mizuk et al. (2019) Yes Adachi et al. (2023)
+Table 6: Presenting theoretical mass and branching ratio predictions, along with assignments for mixed S=0 and S=2 $Y_{c}$ tetraquark states.
+The mass results are given in units of MeV.
+Y(4390) indicate the states observed in $\pi^{+}\pi^{-}h_{c}$ process.
+Y’(4390) indicate the states observed in $\pi^{+}\pi^{-}\psi$ process.
+Mass 4153 4214 4300 4313 4315 4347 4391
+$\omega\chi_{c0}$ 1 54 4 36 99
+Data X(4160) Y(4230) Y(4300) Y(4340) Y(4390)
+Mass 4413 4467 4488 4496 4518 4536 4562
+$\omega\chi_{c0}$ 9 19 15 77 17 51
+Data Y’(4390) Y(4484) Y(4484) Y(4544)
+Mass 4574 4575 4584 4592 4604 4614 4621
+$\omega\chi_{c0}$ 7 131 66 4
 Data Y(4660)
 III.2 Y(4230)
-In the past two decades, the Y(4230) is an exotic particle observed and confirmed by many experimental collaborations, and is considered part of the broader family of charmonium-like states.
-First detected in electron-positron annihilation experiments with a mass of approximately 4.23 GeV Aubert et al.
-(2005), placing it in the region above the open-charm threshold.
+In the past two decades, the Y(4230) is an exotic particle observed and confirmed by many experimental collaborations, and is considered part of the broader family of charmonium-like states Navas et al.
+(2024).
+The Y(4260), the original name of Y(4230), was observed by BaBar in the $\pi^{+}\pi^{-}J/\psi$ invariant-mass spectrum Aubert et al.
+(2005), and was subsequently confirmed by CLEO Collaboration He et al.
+(2006) and Belle Collaboration Yuan et al.
+(2007) in the same process.
+With improved statistical precision, BESIII identified an asymmetry in the cross section, which caused the peak position to move to a lower mass Ablikim et al.
+(2017b).
+The Y(4260) was therefore renamed Y(4230).
 Unlike conventional quarkonium, which consists of a quark-antiquark pair, the Y(4230) is classified as an exotic hadron due to its unconventional properties, including anomalous decay channels and unexpected production mechanisms.
@@
 (2015).
-Y(4230) has been observed by many decay processes but mainly from electron positron collision, where the processes include $e^{+}e^{-}\to\pi^{+}\pi^{-}\psi$Ablikim et al.
-(2020a, 2022b, 2021), $\eta J\psi$Ablikim et al.
-(2020b), $\omega\chi_{c0}$blikim et al.
-(2015), $\pi^{+}\pi^{-}h_{c}$blikim et al.
-(2017), and $K^{+}K^{-}J/\psi$Ablikim et al.
-(2022a), which are listed in Table 5.
-There are some clues about states identification appearing when the resonances are observed from different channels.
-The resonance parameter of Y(4230) in the invariant mass of $\omega\chi_{c0}$blikim et al.
-(2015) is inconsistent with the line shape of the Y(4260) observed in $\pi^{+}\pi^{-}J/\psi$Aubert et al.
-(2005) which is the first observed Y states in 2005 by BABAR.
-This Y(4260) state later superseded by a updated workLees et al.
-(2012) in PDG because two interfering resonances are not excluded.
-But both work were from a single-resonance fit.
-Since the Y(4230) was also observed in $K^{+}K^{-}J/\psi$Ablikim et al.
-(2022a) process, and the resonance parameters is consistent with those from Y(4230) observed from $\pi^{+}\pi^{-}J/\psi$ channel, one may classify the Y(4230) from $K^{+}K^{-}J/\psi$Ablikim et al.
-(2022a) and $\pi^{+}\pi^{-}J/\psi$ to be a same state.
-Later, Y(4230) is observed in the process $\pi^{+}\pi^{-}h_{c}$blikim et al.
-(2017), whose parameters are different from those of Y(4260) observed in $\pi^{+}\pi^{-}J/\psi$ channel in PDG before year 2015.
-However, the resonance parameters of Y(4230) in $\pi^{+}\pi^{-}h_{c}$blikim et al.
-(2017) are consistent with those of the resonance observed in $\omega\chi_{c0}$ Ablikim et al.
+[…]
+[…]
+The Y(4230) was also observed in $K^{+}K^{-}J/\psi$ Ablikim et al.
+(2022a) process, and the resonance parameters were consistent with those from Y(4230) observed from $\pi^{+}\pi^{-}J/\psi$ channel, one may classify the Y(4230) from $K^{+}K^{-}J/\psi$ and $\pi^{+}\pi^{-}J/\psi$ to be a same state.
+Later, the Y(4230) was observed in the process $\pi^{+}\pi^{-}h_{c}$ Ablikim et al.
+(2017a), whose resonance parameters were consistent with those of the resonance observed in $\omega\chi_{c0}$ Ablikim et al.
 (2015), one may support that the Y(4230) from $\omega\chi_{c0}$ and $\pi^{+}\pi^{-}h_{c}$ are likely a same state.
-The discussion above may support to separate Y(4230) into two different states.
-In assignment Table 6, the one from $\pi^{+}\pi^{-}J/\psi$ and $K^{+}K^{-}J/\psi$ might be tentatively named Y(4230) while the one observed from $\omega\chi_{c0}$ and $\pi^{+}\pi^{-}h_{c}$ might be tentatively named Y’(4230) respectively.
-The Y(4230) are assigned to be E1 state of P-wave tetraquark with quantum number $1^{--}$ and Y’(4230) are assigned to be E2 state due to their different decay ratio, where E2 state has much larger ratio for $\omega\chi_{c0}$ channel.
+The number of states in this mass region remains to be further determined experimentally Navas et al.
+(2024).
+This study predicts only one state in the mass region.
+As shown in Table 6, we may assign the Y(4230), observed in the $\pi^{+}\pi^{-}J/\psi$, $K^{+}K^{-}J/\psi$, $\omega\chi_{c0}$, and $\pi^{+}\pi^{-}h_{c}$ channels, to be the second tetraquark state, which is based on its large decay ratio to $\omega\chi_{c0}$, with a corresponding theoretical mass of 4214 MeV.
 III.3 Y(4360)
@@
 This anomalous behavior suggests that Y(4360) could be an exotic hadron.
-As a PDG established state, in PDG data listed in Table 5, the mass of $\psi(4360)$ are from 4.3 to 4.4 GeV, which might show the two different states in the large range.
-One is around 3.4 GeV, and another is around 3.8GeV, which are renamed as Y(4340) and Y(4380) in the assignment respectively.
-The Y(4340) observed only in $\pi^{+}\pi^{-}J/\psi$ process is assigned to be E3 tetraquark state taking a small ratio to $\omega\chi_{c0}$ process which is similar to E1 Y(4230) state, one may be the excitation state in same direction with Y(4230) from $\pi^{+}\pi^{-}J/\psi$ process.
-Similar with the discussion for Y(4230), Y(4380) is also observed in $\pi^{+}\pi^{-}J/\psi$ Ablikim et al.
+As a PDG established state, in PDG data listed in Table 5, the mass of $\psi(4360)$ are from 4.3 to 4.4 GeV, which might show the three different states in the large range.
+The first state is around 4.30 GeV, the second state is around 4.34 GeV, and the third state is around 4.39 GeV, which are renamed as Y(4300), Y(4340), and Y(4390) in the assignment respectively.
+The Y(4300) and Y(4340) observed in $\pi^{+}\pi^{-}J/\psi$ and $\pi^{+}\pi^{-}\psi(2S)$ process are assigned to be S=2 tetraquark state, with a corresponding theoretical mass of 4300 MeV and 4347 MeV respectively.
+The Y(4390) was observed in $\pi^{+}\pi^{-}J/\psi$ Ablikim et al.
 (2021) and $\pi^{+}\pi^{-}h_{c}$ Ablikim et al.
-(2017) respectively, where there line shapes is inconsistent with each other Ablikim et al.
+(2017a) channels respectively.
+In the $\pi^{+}\pi^{-}h_{c}$ process Ablikim et al.
+(2017a), the Y(4230) was also observed.
+The resonance parameters of the Y(4230) in $\pi^{+}\pi^{-}h_{c}$ process were consistent with those of the Y(4230) observed in $\omega\chi_{c0}$ Ablikim et al.
+(2015), which may suggest that the Y(4390) could also be observed in the $\omega\chi_{c0}$ channel.
+In the assignment Table 6, the Y(4390) observed in the $\pi^{+}\pi^{-}h_{c}$ channel is assigned based on its large decay ratio to $\omega\chi_{c0}$, with a corresponding theoretical mass of 4391 MeV.
+The Y(4390) could potentially be observed in the $\omega\chi_{c0}$ channel, which may motivate future experimental searches.
+In Ref Ablikim et al.
+(2017a), the line shapes of Y(4390) observed in $\pi^{+}\pi^{-}h_{c}$ process Ablikim et al.
+(2017a) are inconsistent with those in $\pi^{+}\pi^{-}J/\psi$ and $\pi^{+}\pi^{-}\psi(2S)$ by BaBar Aubert et al.
+(2005, 2007); Lees et al.
+(2014, 2012) and by Belle Yuan et al.
+(2007); Wang et al.
+(2007, 2015); Liu et al.
 (2013).
-One may support that there are two Y states around the mass 4380MeV, which are tentatively named Y(4380) for the one from $\pi^{+}\pi^{-}J/\psi$, and Y’(4380) for the one from $\pi^{+}\pi^{-}h_{c}$.
-In the assignment Table 6, E5 state shows similar property with the Y state from $\omega\chi_{c0}$ and $\pi^{+}\pi^{-}h_{c}$, where E5 state has larger ratio to $\omega\chi_{c0}$ process comparing with E4 state, one may assign Y(4380) and Y’(4380) to be E4 and E5 state respectively.
+One may support that there are two Y states around the mass 4390MeV, which are tentatively named Y(4390) for the one from $\pi^{+}\pi^{-}h_{c}$, and Y’(4390) for the one from $\pi^{+}\pi^{-}J/\psi$.
+In the assignment Table 6, the Y(4390) observed in the $\pi^{+}\pi^{-}\psi$ channel is assigned based on the small decay ratio to $\omega\chi_{c0}$, with a corresponding theoretical mass of 4413 MeV.
 III.4 Y(4660)
@@
 (2015).
-In the assignment Table 6, the Y(4660) is mostly observed in the $\pi^{+}\pi^{-}\psi$ channel, which might be naturally grouped with Y(4230), Y(4340) and Y(4380) from the same process.
-One may assign it to be E15 tetraquark state which has a small decay ratio to $\omega\chi_{c0}$.
+In the assignment Table 6, the Y(4660) is mostly observed in the $\pi^{+}\pi^{-}\psi$ channel, which might be naturally grouped with Y(4230), Y(4300), Y(4340) and Y’(4390) from the same process.
+One may assign it to be tetraquark state which has a small decay ratio to $\omega\chi_{cJ}$, with a corresponding theoretical mass of 4621 MeV.
 III.5 Other $Y_{c}$ states
+The $X(4160)$ was first observed by the Belle Collaboration in double-charmonium production, $e^{+}e^{-}\to J/\psi X(4160)$ with $X(4160)\to D^{+}D^{-}$Pakhlov et al.
+(2008).
+Later, a related state was observed by the LHCb Collaboration in the decay $B^{+}\to J/\psi\phi K^{+}$Aaij et al.
+(2021), reconstructed from the invariant mass spectrum of $J/\psi\phi$.
+The two states observed around 4.15 GeV are grouped together as a single entry in the PDG Navas et al.
+(2024), according to their similar masses and decay widths.
+Whether these observations correspond to the same state still requires further experimental confirmation.
+The nature of the $X(4160)$ therefore remains uncertain and requires further theoretical and experimental investigation.
+The production rate of the $X(4160)$ in double-charmonium processes, along with its decay patterns into the $D^{+}D^{-}$ and $J/\psi\phi$ channels, does not align well with expectations for conventional $c\bar{c}$ states Chao (2008); Navas et al.
+(2024).
+But, some studies have still considered interpretations of the $X(4160)$ as a higher charmonium state Yang et al.
+(2010).
+The $X(4160)$ is more likely interpreted as an exotic hadron, with several theoretical scenarios proposed.
+These include a compact tetraquark configuration, either hidden-charm $c\bar{c}q\bar{q}$ or $c\bar{c}s\bar{s}$, supported by QCD sum rules Chen et al.
+(2017), as well as a meson-meson molecular state, in particular a $D_{s}^{*}\bar{D}_{s}^{*}$ bound state Wang et al.
+(2018).
+Despite these scenarios, the quantum numbers of the $X(4160)$ remain unclear.
+The Belle Collaboration did not measure its quantum numbers Pakhlov et al.
+(2008), while LHCb analyses favor an assignment of $J^{PC}=2^{-+}$ Aaij et al.
+(2021).
+Given its measured mass near 4.16 GeV, the $X(4160)$ may also be interpreted as a tetraquark candidate, based on mass consistency with theoretical predictions for hidden-charm exotics in this region Chen et al.
+(2017).
+However, further experimental confirmation is still required.
+In the assignment, $X(4160)$ is tentatively assigned to be tetraquark state due to masses agreement, with a corresponding theoretical mass of 4153 MeV.
 […]
@@
 III.6 Y(10753)
+Table 7: Theoretical mass, branching ratio predictions, and assignments for mixed S=0 $Y_{b}$ tetraquark states. Mass values are in MeV.
+En E1 E2 E3 E4 E5 E6 E7 E8
+Mass 10497 10542 10573 10685 10749 10769 10790 10808
+$\omega\chi_{b0}$ 1 54 4 36 99 9 19 15
+$\omega\chi_{b1}$ 3 161 12 109 297 27 57 45
+$\omega\chi_{b2}$ 5 269 20 181 496 45 94 75
+Data Y(10753)
 The Y(10753) bottomonium-like state was firstly observed by Belle in the $e^{+}e^{-}\to\pi^{+}\pi^{-}\Upsilon(nS)$ process Mizuk et al.
@@
 The theoretical results for P-wave tetraquarks are compared with selected exotic states, known as Y states, and a tentative assignment is proposed.
-The PDG established state around 4.23 GeV may split into two distinct states.
-In the assignment, the Y(4230), observed in the $\pi^{+}\pi^{-}J/\psi$ and $K^{+}K^{-}J/\psi$ channels, is assigned as E1 state of an S=0 $1^{--}$ P-wave tetraquark.
-The Y’(4230), observed in the $\omega\chi_{c0}$ and $\pi^{+}\pi^{-}h_{c}$ channels, is assigned as E2 state of an S=0 $1^{--}$ P-wave tetraquark.
-The PDG established state around 4.36 GeV may also split into three distinct states.
-In the assignment, the Y(4340) observed only in the $\pi^{+}\pi^{-}J/\psi$ process, is assigned as E3 state of an S=0 $1^{--}$ tetraquark, which can be paired with the E1 Y(4230) state.
-The Y(4380), observed in the $\pi^{+}\pi^{-}J/\psi$ channel, is assigned as E4 state of an S=0 $1^{--}$ tetraquark, while the Y’(4380), observed in the $\pi^{+}\pi^{-}h_{c}$ channel, is assigned as E5 state of an S=0 $1^{--}$ tetraquark.
-The PDG established state Y(4660), observed in the $\pi^{+}\pi^{-}J/\psi$ channel and grouped with Y(4230), Y(4340) and Y(4380) from the same process, is assigned as the E15 S=0 $1^{--}$ tetraquark state.
-The Y(4484), observed in the $K^{+}K^{-}J/\psi$ channel by BESIII, is grouped with Y(4230) from the $\pi^{+}\pi^{-}J/\psi$ channel and tentatively assigned as the E8 S=0 $1^{--}$ tetraquark state.
-The Y(4544), newly observed in the $\omega\chi_{c1}$ channel by BESIII, is grouped with Y’(4230) from the $\omega\chi_{c0}$ channel and tentatively assigned as E11 S=0 $1^{--}$ tetraquark state.
+The PDG established state around 4.23 GeV may be treated as a single state at this stage.
+In the assignment, the Y(4230), observed in the $\pi^{+}\pi^{-}J/\psi$, $K^{+}K^{-}J/\psi$, $\omega\chi_{c0}$ and $\pi^{+}\pi^{-}h_{c}$ channels, is assigned as S=0 $1^{--}$ tetraquark state.
+The PDG established state around 4.36 GeV may split into four distinct states.
+In the assignment, the Y(4300) and Y(4340) observed only in the $\pi^{+}\pi^{-}J/\psi$ process, are assigned as $1^{--}$ tetraquark state respectively.
+The Y(4390), observed in the $\pi^{+}\pi^{-}h_{c}$ channel, is assigned as a $1^{--}$ tetraquark state with a large decay ratio to $\omega\chi_{cJ}$, while the Y’(4390), observed in the $\pi^{+}\pi^{-}J/\psi$ channel, is assigned as a tetraquark state with a small decay ratio to $\omega\chi_{cJ}$.
+The Y(4390) could potentially be observed in the $\omega\chi_{c0}$ channel, which may motivate future experimental searches.
+The PDG established state Y(4660), observed in the $\pi^{+}\pi^{-}J/\psi$ channel and grouped with Y’(4390) from the same process, is assigned as the a $1^{--}$ tetraquark state with a small decay ratio to $\omega\chi_{cJ}$.
+The Y(4484), observed in the $K^{+}K^{-}J/\psi$ channel by BESIII, and the Y(4544), newly observed in the $\omega\chi_{c1}$ channel by BESIII, are tentatively assigned as $1^{--}$ tetraquark states due to mass agreement, respectively.
 The Y(10753), observed in $\pi^{+}\pi^{-}\Upsilon(nS)$ and $\omega\chi_{bJ}(1P)$ channel by Belle, is assigned as an E5 S=0 $1^{--}$ bottomonium-like tetraquark state.
@@
 $\psi_{NL}$ $=\sum_{{n_{\chi_{i}},l_{\chi_{i}}}}A(n_{\chi_{1}},n_{\chi_{2}},n_{\chi_{3}},l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}})$ (12)
-×$\times\psi_{n_{\chi_{1}}l_{\chi_{1}}}(\vec{\chi}_{1}\,)\otimes\psi_{n_{\chi_{2}}l_{\chi_{2}}}(\vec{\chi}_{2}\,)\otimes\psi_{n_{\chi_{3}}l_{\chi_{3}}}(\vec{\chi}_{3})$
+$\times\psi_{n_{\chi_{1}}l_{\chi_{1}}}(\vec{\chi}_{1}\,)\otimes\psi_{n_{\chi_{2}}l_{\chi_{2}}}(\vec{\chi}_{2}\,)\otimes\psi_{n_{\chi_{3}}l_{\chi_{3}}}(\vec{\chi}_{3})$
 where $\psi_{n_{\chi_{i}}l_{\chi_{i}}}$ are harmonic oscillator wave functions.
 The sum $n_{\chi_{i}},l_{\chi_{i}}$} is over $n_{\chi_{1}},n_{\chi_{2}},n_{\chi_{3}},l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}$.
-$N$ and $L$ are respectively the total principle quantum number and total angular momentum, and $L=l_{\chi_{1}}+l_{\chi_{2}}+l_{\chi_{3}}=1$ for P wave tetraquark and $N=(2n_{\chi_{1}}+l_{\chi_{1}})+(2n_{\chi_{2}}+l_{\chi_{2}})+(2n_{\chi_{3}}+l_{\chi_{3}})$.
+$N$ and $L$ are respectively the total principle quantum number and total angular momentum, and $L=l_{\chi_{1}}+l_{\chi_{2}}+l_{\chi_{3}}=1$ while $l_{\chi_{1}},\,l_{\chi_{2}},\,l_{\chi_{3}}\leq 1$ for P wave tetraquark and $N=(2n_{\chi_{1}}+l_{\chi_{1}})+(2n_{\chi_{2}}+l_{\chi_{2}})+(2n_{\chi_{3}}+l_{\chi_{3}})$.
 The complete bases of the tetraquarks are listed in Table 8 up to $N=3$ for reference.
@@
 $\psi_{0,1}(\vec{\chi}_{1}\,)\psi_{0,0}(\vec{\chi}_{2}\,)\psi_{1,0}(\vec{\chi}_{3}\,)$
-Appendix B Cross term between different configuration from spin-spin interaction
+Appendix B Cross term between different configuration from central potential
 Table 9: Expectation values of the color-spin operator for the S=0 tetraquark state.
-$<\psi^{k1}_{c}\psi^{h1}_{s}|\vec{\lambda}_{i}\vec{\lambda}_{j}\vec{\sigma}_{i}\vec{\sigma}_{j}|\psi^{k2}_{c}\psi^{h2}_{s}>$
+$\psi^{k1}_{c}\psi^{h1}_{s}|\vec{\lambda}_{i}\vec{\lambda}_{j}\vec{\sigma}_{i}\vec{\sigma}_{j}|\psi^{k2}_{c}\psi^{h2}_{s}>$
 $\vec{\lambda}_{i}\vec{\lambda}_{j}\vec{\sigma}_{i}\vec{\sigma}_{j}$ $|\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(0\otimes 0)}\rangle$ $|\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(1\otimes 1)}\rangle$ $|\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(0\otimes 0)}\rangle$ $|\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(1\otimes 1)}\rangle$
@@
 Table 10: Expectation values of the color-spin operator for S=2 tetraquark state.
-$<\psi^{k1}_{c}\psi^{h1}_{s}|\vec{\lambda}_{i}\vec{\lambda}_{j}\vec{\sigma}_{i}\vec{\sigma}_{j}|\psi^{k2}_{c}\psi^{h2}_{s}>$
+$\psi^{k1}_{c}\psi^{h1}_{s}|\vec{\lambda}_{i}\vec{\lambda}_{j}\vec{\sigma}_{i}\vec{\sigma}_{j}|\psi^{k2}_{c}\psi^{h2}_{s}>$
 $\vec{\lambda}_{i}\vec{\lambda}_{j}\vec{\sigma}_{i}\vec{\sigma}_{j}$ $|\psi^{c}_{\bar{3}\otimes 3}\psi^{S=2}_{(1\otimes 1)}\rangle$ $|\psi^{c}_{6\otimes\bar{6}}\psi^{S=2}_{(1\otimes 1)}\rangle$
@@
 $|\psi^{c}_{6\otimes\bar{6}}\psi^{S=2}_{(1\otimes 1)}\rangle$ $\left(0,4\sqrt{2},-4\sqrt{2},-4\sqrt{2},4\sqrt{2},0\right)$ $\left(\frac{4}{3},-\frac{10}{3},-\frac{10}{3},-\frac{10}{3},-\frac{10}{3},\frac{4}{3}\right)$
-The color-spin operator $\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j}$ in the Hamiltonian in Eq.
-(II.2), along with the mass-dependent factor in the spin-spin interaction,
+The color operator $\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}$ in the Hamiltonian in Eq.
+(II.2), along with the mass-dependent coefficient in the Cornell-like potential, mixes different color configurations, leading to nonzero cross terms for charmonium-like and bottomonium-like tetraquarks.
+Similarly, the color-spin operator $\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j}$, along with the mass-dependent factor in the spin-spin interaction,
 $\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}V_{ss}$ $=\frac{2B_{ij}\sigma_{ij}^{3}}{3m_{i}m_{j}\sqrt{\pi}}e^{-\sigma_{ij}^{2}r_{ij}^{2}}\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j},$ (13)
-mixes different color-spin configurations, leading to nonzero cross terms for charmonium-like and bottomonium-like tetraquarks.
+mixes different color-spin configurations, also leading to nonzero cross terms for charmonium-like and bottomonium-like tetraquarks.
 For reference, the expectation values of $O_{ij}=\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j}$ for S=0 and S=2 states are listed in Table 9 and Table 10 respectively, with all components given in the format $(O_{12},O_{13},O_{23},O_{14},O_{24},O_{34})$.